2. In a Math test, the mean score is 45 and the standard deviation is 4. Assuming normality, what is the probability


that a score picked at random will lie


ve score 50?


b. below score 38?


a.


3.


Consider the normal distribution of IQs with a mean of 100 and a standard dev

1
Expert's answer
2022-04-11T14:57:35-0400

"P(X>50)=1-P(X<50)=1-P(Z<\frac{x-\mu}{\sigma})=1-P(Z<\frac{50-45}{4})=1-(Z<1.25)=1-0.89435=0.10565"

b. "P(X<38)=P(Z<\frac{x-\mu}{\sigma})=P(Z<\frac{38-45}{4})=P(Z<-1.75)=0.04006"